# A semi-supervised framework for diverse multiple hypothesis testing scenarios

### v2

@@ 7 Discussion
We used the exact same parameters in the vignettes https://github.com/martinjzhang/AdaFDRpaper/tree/master/vignettes to implement AdaFDR on the 10 datasets that Zhang et al. look at. Interestingly, the two fMRI datasets only consider the Brodmann label as the side information. We tried using all the available side information, but the statistical power was worse (data not shown). In the gene-drug response data, we used AdaFDR with fast_mode = False. This is in accordance to the discussion from their paper that the fast version of AdaFDR is recommended when few discoveries are expected or if the number of hypotheses is small (which is not the case for both of the gene-drug response datasets).
S.11 Computer specifications for computation times
@@
S.12 Data preparation and searching with HEK293 data
We downloaded the human proteome (UP000005640, downloaded on 2023/09/23 from UniProt) and prepared a target peptide database along with 5 randomly shuffled decoy peptide databases using Tide-index within the Crux Toolkit v4.1.6809338 [43, 29] with all options set to default. An output containing the target and decoy peptide pairs are conveniently provided using the --peptide-list T option in Tide-index. Each of the RAW 24 HEK293 spectrum files [9] were converted to mzML format using MSConvert 3.0.22314 with the vendor peak-picking filter using the default settings. For each of the 5 decoy databases, we searched each spectrum from the combined 24 mzML spectrum files against the combined target-decoy peptide database using Tide-search [15], using the options --top-match 1 --auto-precursor-window warn --auto-mz-bin-width warn --concat T. The resulting search files were then converted to so-called pin files using the make-pin function in Crux. Each row in the pin files correspond to the optimal peptide-spectrum match (PSM) for each spectrum, the primary score for quantifying this match, called XCorr scores, and a collection of auxiliary information regarding the PSM.
@@
Table S4: List of described side information used by HEK293 data. The list of side information used by RESET and Adaptive Knockoffs in the HEK293 data, adapted from https://crux.ms/file-formats/features.html. Note that the Charge state is represented as a one-hot vector, where we left out a Charge state of +1 (due to linearity, since the one-hot vector sums to one).
Search algorithms like the one above are typically ‘spectrum-centric’, meaning that for each spectrum, we look for the best matching database peptide, as oppose to the other way round. Consequently, the same database peptide may be matched multiple times with different scores, or have no match at all. Hence for each target peptide, we define $Z_{i}$ as the maximum XCorr score of all the PSMs associated to that target peptide. Similarly for each decoy peptide, we define $\tilde{Z}_{i}$ as the maximum XCorr score of all the PSMs associated to that decoy peptide. For target or decoy peptides that do not appear in the pin file, they get a score of $\infty$- ∞. Then we compete each target and its corresponding paired decoy by recording the winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ along with the label $L_{i}=sign(Z_{i}-\tilde{Z}_{i})$, randomly breaking any ties. Peptides with a winning score of $\infty$- ∞, indicating that neither the target or its paired decoy are in the pin file, are thrown out. We assign each winning peptide the side information defined as the auxiliary information $\mathbf{x}_{i}$ of the underlying PSM that had the XCorr score of $W_{i}$. The side information that was subsequently used by RESET and Adaptive Knockoffs is given in Table S4.
S.13 Estimating the computation times of Adaptive Knockoffs
Running Adaptive Knockoffs (AdaKO) on all peptide datasets would be infeasible. Therefore, we resorted to estimating its runtimes by considering the time it took to complete a single iteration that determines which hypothesis to reveal next, and multiplying this value by the estimated number of iterations before the procedure stops.
Since the time it takes a single iteration depends on the number of hypotheses that were already revealed, we took different such measurements as we varied this number, and averaged them. Specifically, as AdaKO does, we first revealed the labels that were scored less than or equal to the 10% quantile of non-zero scoring hypotheses, and recorded the time it takes for a single interation from this point. To be clear, the revealed labels thus includes those zero-scoring hypotheses since they will be less than this score cutoff — it is only the 10% quantile that considers the non-zero scoring hypotheses. Then we took similar measurements by increasing the ‘10%’ value, in 10% increments, up to a value of ‘10k%’ where we define $k$ next.
We chose the value of $k$ so that the proportion of revealed labels, $10k\%$, estimates the number of revealed labels when AdaKO terminates at the 1% FDR level. That way, the $k$ times should provide an accurate sample of the number single iteration times across all of AdaKO’s iterations until it terminates. Of course, we are unable to compute when Adaptive Knockoffs might terminate, so we estimated this using the value $q$, the number of target and decoy peptides that scored above the 1% FDR cutoff using RESET Ensemble. Then we defined $k=\lfloor(m-q)/m\cdot 10\rfloor$, where $m$ is the total number of hypotheses (peptides) considered by Adaptive Knockoffs. We estimated the runtime by
$\bar{t}_{k}\cdot(m^{\prime}-q),$
where $\bar{t}_{k}$ is the average of the $k$ measured single-iteration times, and $m^{\prime}$ is the number of labels that are yet to be revealed at the start of AdaKO’s iterations, i.e., the number of labels with scores above the 10% quantile of non-zero scoring hypotheses. Thus the value $m^{\prime}-q$ is our estimate of the number of iterations until Adaptive Knockoffs stop.
In Section 5.2.2, we considered 10 repeated applications of RESET Ensemble, varying RESET Ensemble’s internal seed. In this case, we define $\hat{\ell}$ as the average of the number of winning target and decoy peptides above the cutoff over each application.
S.14 Data preparation and searching with PRIDE-20 data
@@
We followed the preparation and searching outlined in [18]. The only difference is that we used an updated version of the Crux toolkit. We provide the following brief description for the reader’s convenience. Twenty spectrum files and protein databases from separate projects, referred to as the PRIDE-20 dataset, were downloaded. Tide-index was used to prepare the target peptide database along with 10 randomly shuffled decoy peptide databases for each of the 20 spectrum files. Tide-search was used to search each spectrum file against the 10 combined target-decoy databases using two modes: ‘narrow’ and ‘open’ search options as they outline. Briefly a ‘narrow’ search, which is the default search mode, only searches each spectrum in the spectrum file against peptides in the target-decoy database that have a theoretical mass within a small tolerance of the sample peptide’s mass that generated the experimental spectrum. On the other hand, an ‘open’ search makes this mass-tolerance larger so that a spectrum may match with a peptide with a completely different mass. This is desirable at times, since sample peptides can undergo post-translational modifications which modify the mass of the peptide and therefore the experimental spectrum. Hence when searching in an ‘open’ mode, the modified spectrum may still be correctly matched. The search files were then converted to pin files using the make-pin function in Crux and subsequently filtered for the top 1 (the optimal) peptide-spectrum match for each spectrum.
@@
For each spectrum file and search type, we obtained the triples $(L,W,\mathbf{x})$ in the following way. For each spectrum file that had no modifications (we explain this in more detail next), we proceeded essentially in the same way as in Section S.12. That is, for each peptide in the pin file, we define $Z_{i}$ (if a target peptide) or $\tilde{Z_{i}}$ (if a decoy peptide) as the maximum Tailor score of all the PSMs associated to the target peptide [47]. Here we are using the more sensitive Tailor scores and relegate the XCorr scores as one of the side information variables. Then we compete each target and its corresponding paired decoy in the same way as Section S.12, to obtain the triple $(L_{i},W_{i},\mathbf{x}_{i})$ where the side information $\mathbf{x}_{i}$ is obtained from the auxiliary information of the underlying PSM that had the Tailor score of $W_{i}$. A description of the side information used is given in Table S5.
@@
Table S5: List of described side information used by PRIDE-20 data. The list of side information used by RESET and Adaptive Knockoffs in the PRIDE-20 data, adapted from https://crux.ms/file-formats/features.html. Note that the Charge state is represented as a one-hot vector, where we left out a Charge state of +1 (due to linearity, since the one-hot vector sums to one). XCorr is used as side information since we define $W$ in terms of Tailor scores here instead.
@@
We next consider spectrum files that were searched with variable modifications (---auto-modifications T). Here we need to make a slight adjustment to account for some dependency within the data. Specifically, using variable modifications creates several ‘copies’ of the peptides in the target-decoy database that are distinguished only by slight alterations to the mass of some of the amino acids. As an example, PEPTIDE may generate the following ‘copies’: PE[16]PTIDE, PEPTIDE[16] and PE[16]PTIDE[16], where [16] indicates an increase of 16 Daltons to the amino acid on the left. Consequently, each of these peptides are usually correlated in the data — i.e., if one of these peptides is scored high, than usually they all are. Hence in order to satisfy Assumption 3, we follow the protocol outlined in [18]. That is, we identify all the copies as being the same peptide, ‘PEPTIDE’, with each of the variable modifications ignored. Then we proceed with the same steps as the above paragraph, by determining the maximum score associated with PEPTIDE and recording it as $Z_{i}$ (or $\tilde{Z}_{i}$ if it is a decoy). We define the labels, winning scores and side information in the same way to obtain our triplets $(L,W,\mathbf{x}_{i})$.
@@
Due to runtime considerations, we used 13 of the smallest narrow search files to assess Adaptive Knockoff’s power as described in Section 5.2.2 of the manuscript, and only one of the 10 target-decoy databases. To fairly compare with Adaptive Knockoffs, we used the same target-decoy databases for RESET ensemble. All 20 spectrum files and all 10 target-decoy databases for each spectrum file were used in the comparison of RESET ensemble and FDP-SD while controlling the FDP.
S.15 Brief justification of the assumptions used with HEK293 and PRIDE-20 datasets
A standard assumption of the mass spectrometry literature, which is also empirically validated, is that the label $L_{i}$, indicating whether an incorrect target or its decoy scored higher, is uniformly ±$\pm 1$ independently of all other target-decoy pairs and scores [21, 39]. RESET additionally requires this statement to be true if we further condition on the side information. This is obviously true for some side information variables which are constant between the target and corresponding decoys — e.g., PepLen, Charge, lnNumSP, dm, absdM. The other side information variables are score-related — e.g., Xcorr, deltCn and deltLCn, and so conditioning on them assumes no more than the ‘standard assumption’. Hence, we believe Assumption 3 with $P(L_{i}=1)=c_{0}=1/2$ to be reasonably satisfied.
S.16 Preparation of data with p-values

### v3

@@ C.11 Further implementation details of AdaFDR
We used the exact same parameters in the vignettes https://github.com/martinjzhang/AdaFDRpaper/tree/master/vignettes to implement AdaFDR on the 10 datasets that Zhang et al. look at. Interestingly, the two fMRI datasets only consider the Brodmann label as the side information. We tried using all the available side information, but the statistical power was worse (data not shown). In the gene-drug response data, we used AdaFDR with fast_mode = False. This is in accordance to the discussion from their paper that the fast version of AdaFDR is recommended when few discoveries are expected or if the number of hypotheses is small (which is not the case for both of the gene-drug response datasets). For the same reasons, we used AdaFDR with fast_mode = True on Simulation 5 when the number of hypotheses was small, and fast_mode = False on Simulations 7 and 8 when the number of hypotheses were large.
C.12 Simulation details for p-value setting
C.12.1 Simulation 6
Simulation 6 in Section 6.2.2 models “noisy” side information since only the first 2 of the 100 draws are used to distinguish between the true and false null hypotheses. Each p-value $p_{i}$ is drawn i.i.d. from a two-group beta mixture model, where the true nulls have a uniform distribution and the false nulls have an alternative beta distribution. Specifically, the beta mixture model is given by:
where $\mathbf{x}_{i}\in[0,1]^{100}$ denotes the side information of the $i$th hypothesis, $\pi(\mathbf{x}_{i})$ denotes proportion of the false nulls in the mixture, $h$ denotes the false null distribution of $\text{Beta}(1/\mu(\mathbf{x}_{i}),1)$. The $\pi(\mathbf{x}_{i})$’s and $\mu(\mathbf{x}_{i})$’s are determined by the following relationships with the 100-dimensional side information:
where $\theta=(3,3,0,\dots,0)\in\mathbb{R}^{100}$, $\beta=(2,2,0,\dots,0)\in\mathbb{R}^{100}$ and $\theta_{0}$ is chosen to satisfy $\frac{1}{m}\sum_{i}\pi(\mathbf{x}_{i})=0.3$. Clearly, only the first two values in $\mathbf{x}_{i}\in[0,1]^{100}$ contribute to the beta mixture model, as intended. We generated 100 independent datasets, where in each we drew the 2000 p-values as described.
C.12.2 Simulation 7 (false null probability)
The $i$th hypothesis is randomly identified as a false null dependent on its side information $\mathbf{x}_{i})$ with probability (Figure C11)
$\pi_{i}:=\pi(\mathbf{x}_{i}):=0.1\cdot\sum_{j=0}^{2}w_{j}\cdot f_{j}(\mathbf{x}_{i}),$
where $(w_{0},w_{1},w_{2})=(0.5,0.25,0.25)$, $f_{1}$ and $f_{2}$ are normal densities with parameters $(\mu_{1},\sigma_{1})=(0.25,0.05)$ and $(\mu_{2},\sigma_{2})=(0.75,0.05)$ respectively, truncated between 0 and 1 and normalized such that $\int_{[0,1]}f_{1}dx=\int_{[0,1]}f_{2}dx=1$, and $f_{0}(x):=a\cdot\exp(ax)/(\exp(a)-1)$ where $a=0.5$.
Figure C11: The probabilities of a hypothesis being a false null, $\pi$, as a function of the hypothesis’ side information, $\mathbf{x}$.
C.13 Computer specifications for computation times
@@
C.14 Estimating the computation times of Adaptive Knockoffs
Running Adaptive Knockoffs (AdaKO) on each of the HEK293 and PRIDE-20 spectrum files is impractical. We therefore resorted to estimate its runtime, which requires dealing with two challenges. The first is that the runtime of each iteration of AdaKO, where it determines which hypothesis to reveal next, varies with the number of remaining hypotheses. The second is that we cannot precisely predict the number of iterations AdaKO will take.
Addressing the second challenge first, we note that AdaKO starts by revealing the label of any hypothesis $H_{i}$ for which $W_{i}\leq q^{*}_{0.1}$, where $q^{*}_{c}$ is the $c$-quantile of all positive scoring hypotheses. Let $m^{\prime}$ be the number of remaining hypotheses after this initial reveal. The number of iterations that AdaKO will make before stopping is then $m^{\prime}$-$\ell$, where $\ell$ is the number of hypotheses (target or decoy peptides) that remain when it stops. We can estimate $\ell$ with $\hat{\ell}$, the number of target and decoy peptides that score above the 1% FDR cutoff using RESET Ensemble. Thus we estimate the number of AdaKO iterations by $m^{\prime}$-$\hat{\ell}$.
To estimate the runtime of each iteration we resorted to taking $k$ measurements. Indeed, for each $i\in\{1,\dots,k\}$, we revealed the labels of any hypotheses $H_{i}$ for which $W_{i}\leq q_{(0.1\cdot i)}^{*}$ and determined the time it took for AdaKO to reveal the next label. We chose $k$ to be the largest positive integer such that the number of remaining labels, after revealing the labels $L_{i}$ for which $W_{i}\leq q_{(0.1\cdot k)}^{*}$, is still more than $\hat{\ell}$. Thus, we obtain the runtime of $k$ single iterations that are evenly spaced, starting from AdaKO’s first iteration and ending at approximately AdaKO’s last iteration before it terminates with $\hat{\ell}$ remaining hypotheses. Next, we computed $\bar{t}_{k}$, the average of those $k$ measurements. Thus our overall estimate is:
$\bar{t}_{k}\cdot(m^{\prime}-\hat{\ell}).$
In the analysis of the PRIDE-20 data in Section E.2, we considered 10 repeated applications of RESET Ensemble, varying RESET Ensemble’s internal seed. In this case, we define $\hat{\ell}$ as the average of the number of winning target and decoy peptides above the cutoff over each application.
C.15 Method-related figures
Figure C12: Location of false null hypotheses according to the pair of values $\left(\mathbf{x}_{i1},\mathbf{x}_{i2}\right)$ in the square $\left[-100,100\right]\times\left[-100,100\right]$. The same figure is given in Lei and Fithian (2018).
Appendix D Data preparation
D.1 Data preparation and searching with HEK293 data
We downloaded the human proteome (UP000005640, downloaded on 2023/09/23 from UniProt) and prepared a target peptide database along with 5 randomly shuffled decoy peptide databases using Tide-index within the Crux Toolkit v4.1.6809338 (Park et al. (2008); Kertesz-Farkas et al. (2023)) with all options set to default. An output containing the target and decoy peptide pairs are conveniently provided using the --peptide-list T option in Tide-index. Each of the RAW 24 HEK293 spectrum files (Chick et al. (2015)) were converted to mzML format using MSConvert 3.0.22314 with the vendor peak-picking filter using the default settings. For each of the 5 decoy databases, we performed a ‘narrow’ search on each spectrum from the combined 24 mzML spectrum files against the combined target-decoy peptide database using Tide-search (Diament and Noble (2011)), using the options --top-match 1 --auto-precursor-window warn --auto-mz-bin-width warn --concat T. The resulting search files were then converted to so-called pin files using the make-pin function in Crux. Each row in the pin files correspond to the optimal peptide-spectrum match (PSM) for each spectrum, the primary score for quantifying this match, called XCorr scores, and a collection of auxiliary information regarding the PSM.
@@ D.1 Data preparation and searching with HEK293 data
Table D3: List of described side information used by HEK293 data. The list of side information used by RESET and Adaptive Knockoffs in the HEK293 data, adapted from https://crux.ms/file-formats/features.html. Note that the Charge state is represented as a one-hot vector, where we left out a Charge state of +1 (due to linearity, since the one-hot vector sums to one).
We constructed the target and decoy peptide scores, $Z_{i}$ and $\tilde{Z}_{i}$ respectively, according to our description in Section C.6 using the XCorr score as the PSM scoring function. Subsequently we record the winning score $W_{i}=Z_{i}\vee\tilde{Z}_{i}$ along with the label $L_{i}=sign(Z_{i}-\tilde{Z}_{i})$. In the case of ties ($Z_{i}=\tilde{Z}_{i}$), we randomly set $L_{i}$ to be +1+1 or 1-1 with equal probability. To each pair $(W_{i},L_{i})$ we attribute the side information defined as the auxiliary information $\mathbf{x}_{i}$ of the underlying PSM that had the XCorr score of $W_{i}$. In the case there is more than one PSM with the same XCorr score of $W_{i}$, we randomly select one of the target PSMs if $L_{i}=+1$ or one of the decoy PSMs if $L_{i}=-1$. The side information that was subsequently used by RESET and Adaptive Knockoffs is given in Table D3.
D.2 Data preparation and searching with PRIDE-20 data
@@
We followed the preparation and searching outlined in Freestone et al. (2024). The only difference is that we used an updated version of the Crux toolkit. We provide the following brief description for the reader’s convenience. Twenty spectrum files and protein databases from separate projects, referred to as the PRIDE-20 dataset, were downloaded (see Table 2 in Freestone et al. (2024)). Tide-index was used to prepare the target peptide database along with 10 randomly shuffled decoy peptide databases for each of the 20 spectrum files. Tide-search was used to search each spectrum file against the 10 combined target-decoy databases using two modes: ‘narrow’ and ‘open’ search options as outlined in Freestone et al. (2024). The search files were then converted to pin files using the make-pin function in Crux and subsequently filtered for the top 1 (the optimal) peptide-spectrum match for each spectrum.
@@
For each spectrum file and search type, we obtained the triples $(L,W,\mathbf{x})$ in the following way. For each spectrum file that had no modifications (we explain this in more detail next), we proceeded essentially in the same way as in Section D.1. That is, for each peptide in the pin file, we define $Z_{i}$ (if a target peptide) or $\tilde{Z_{i}}$ (if a decoy peptide) as the maximum Tailor score of all the PSMs associated to the target peptide (Sulimov and Kertész-Farkas (2020)). Here we are using the more sensitive Tailor score and relegate the XCorr score as one of the side information variables. Then we compete each target and its corresponding paired decoy in the same way as Section D.1, to obtain the triple $(L_{i},W_{i},\mathbf{x}_{i})$ where the side information $\mathbf{x}_{i}$ is obtained from the auxiliary information of the underlying PSM that had the Tailor score of $W_{i}$. A description of the side information used is given in Table D4.
@@
Table D4: List of described side information used by PRIDE-20 data. The list of side information used by RESET and Adaptive Knockoffs in the PRIDE-20 data, adapted from https://crux.ms/file-formats/features.html. Note that the Charge state is represented as a one-hot vector, where we left out a Charge state of +1 (due to linearity, since the one-hot vector sums to one). XCorr is used as side information since we define $W$ in terms of Tailor scores here instead.
@@
We next consider spectrum files that were searched with variable modifications (---auto-modifications T). Here we need to make a slight adjustment to account for some dependency within the data. Specifically, using variable modifications creates several ‘copies’ of the peptides in the target-decoy database that are distinguished only by slight alterations to the mass of some of the amino acids. As an example, PEPTIDE may generate the following ‘copies’: PE[16]PTIDE, PEPTIDE[16] and PE[16]PTIDE[16], where [16] indicates an increase of 16 Daltons to the amino acid on the left. Consequently, the scores of all these peptides are usually correlated: if one of these is scored high, than often they all are. Hence in order to satisfy Assumption 3, we follow the protocol outlined in Freestone et al. (2024). That is, we identify all the copies as being the same peptide, ‘PEPTIDE’, with each of the variable modifications ignored. Then we proceed with the same steps as the above paragraph, by determining the maximum score associated with PEPTIDE and recording it as $Z_{i}$ (or $\tilde{Z}_{i}$ if it is a decoy). We define the labels, winning scores and side information in the same way to obtain our triplets $(L,W,\mathbf{x})$.
@@
Due to runtime considerations, we used 13 of the spectrum files to assess Adaptive Knockoff’s power as described in Section 5.2.2 of the manuscript (each resulting in <33K<33K PSMs), and only one of the 10 target-decoy databases. To fairly compare with Adaptive Knockoffs, we used the same target-decoy databases for RESET ensemble. All 20 spectrum files and all 10 target-decoy databases for each spectrum file were used in comparing RESET Ensemble’s FDX control with FDP-SD’s.
D.3 Description of real data with p-values
We provide the following description of each dataset used in Section 6.4 in the main text and in Section F.3. The following three RNA-seq datasets, the proteomics dataset, the two microbiome datasets, the two fMRI datasets, and the two eQTL datasets were downloaded from https://github.com/patrickrchao/AdaPTGMM_Experiments/tree/main/AdaFDR_experiments/data_files. The last gene-drug response data was obtained from the adaptMT package.
  • Three RNA-seq datasets (Airway (Himes et al. (2014)), Bottomly (Bottomly et al. (2011)), Pasilla (Brooks et al. (2011))): The goal is to identify genes that are associated with varying gene expression levels in response to a change of conditions. In the Airway data, the differential expression analysis is conducted in response to a drug, dexamethasone; in the Bottomly data, the analysis is conducted between two mouse strains; and in the Pasilla data, the analysis is conducted between normal and so-called ‘Pasilla knockdown’ conditions (Huber and Reyes ). In all cases the side information consists of the log-normalized gene expression read counts.
  • A proteomics dataset (Proteomics (Dephoure and Gygi (2012))): The goal is to identify proteins that have different protein abundances when treated with rapamycin versus DMSO in yeast cells. Each protein consists of side information equal to the natural log of the number of peptides belonging to the protein that were ‘quantified’ in all six samples.
  • Two microbiome datasets (Microbiome Enigma Ph and Microbiome Enigma Al (Smith et al. (2015); Korthauer et al. (2019))): The goal is to identify microbiome organisms, recorded as OTUs (Operational Taxonomic Units), that are associated with certain conditions (pH and Al). Each hypothesis (OTU) is equipped with a two-dimensional side information — the ubiquity, which is defined as the percentage of samples detected with the OTU, and the mean nonzero abundance, which is defined as the average abundance of each OTU across the samples in which it was detected.
  • Two fMRI datasets (Auditory and Imagination (Tabelow and Polzehl (2011))): The goal is to identify voxels that are activated in response to different stimuli. In the first dataset, a single individual received auditory stimulus and in the second dataset, a single individual was instructed to imagine playing tennis. Each hypothesis (voxel) is equipped with four dimensional side information: the spatial position of the voxel (in three dimensions) and a categorical variable with the Brodmann label (Brodmann (1909)), which is used to delineate different areas of the brain with different functions.
  • Two eQTL datasets (Subcutaneous and Omentum (Consortium et al. (2015); Consortium (2020))): The goal is to identify single nucleotide polymorphisms (SNPs) that are associated with changes in gene expression levels. Such an association is referred to as an expression quantitative trait locus (eQTL). Each hypothesis (SNP) is equipped with the following four side information variables: (1) the distance between the SNP and the gene transcription’s start site (TSS), (2) the log gene expression level, (3) the alternative allele frequency (AAF) of the SNP and (4) the chromatin state of the SNP.
  • Gene-drug response data (Estrogen (Davis and Meltzer (2007); Li and Barber (2017))): The goal is to identify genes that are differentially expressed in response to a low-dosage estrogen treatment applied to breast cancer cells. Each hypothesis (gene) is assigned a ‘rank’ that was generated using an analysis of the gene expression levels in response to (1) a high-dosage estrogen treatment and (2) a medium-dosage estrogen treatment. Here, the ranks are in terms of the strength of association between the expression level of each gene and the level of dosage in (1) and (2). Thus, we can analyse the data twice, once for each side information variable based on (1) and (2).
D.4 Preparation of data with p-values